OpenStudy (anonymous):

Subtract: -18 – (-9) = -18 - +9 = -9 right

OpenStudy (blurbendy):

the top would simplify to be - 18 + 9 = -9

OpenStudy (phi):

two minus signs "turn into" a plus sign so -18 – (-9) = becomes -18 + 9 the - -9 becomes +9 of course -18 + 9 = -9 you said -18 - +9 = -9 right which is confusing... that second - sign should not be there. It should only be -18 + 9 which does equal -9